OpenStudy (anonymous):

find the x and y intercepts: -2y+x=2y-8

OpenStudy (bahrom7893):

x intercept => y = 0 0 + x = 0 - 8 x = -8, so: (-8;0)

OpenStudy (bahrom7893):

y intercept => x = 0 -2y + 0 = 2y - 8 -2y-2y = -8 -4y = -8 y = 2

OpenStudy (bahrom7893):

thus y intercept: (0;2)